# FOSP: Fine-tuning Offline Safe Policy through World Models

### v1

@@ Abstract
Model-based Reinforcement Learning (RL) has shown its high training efficiency and capability of handling high-dimensional tasks. Regarding safety issues, safe model-based RL can achieve nearly zero-cost performance and effectively manage the trade-off between performance and safety. Nevertheless, prior works still pose safety challenges due to the online exploration in real-world deployment. To address this, some offline RL methods have emerged as solutions, which learn from a static dataset in a safe way by avoiding interactions with the environment. In this paper, we aim to further enhance safety during the deployment stage for vision-based robotic tasks by fine-tuning an offline-trained policy. We incorporate in-sample optimization, model-based policy expansion, and reachability guidance to construct a safe offline-to-online framework. Moreover, our method proves to improve the generalization of offline policy in unseen safety-constrained scenarios. Finally, the efficiency of our method is validated on simulation benchmarks with five vision-only tasks and a real robot by solving some deployment problems using limited data.
Keywords: Safe Model-based Reinforcement Learning, Offline-to-Online
@@ 1 Introduction
Model-based Reinforcement Learning (MBRL) has proven to be a powerful tool to enhance sample efficiency by constructing a world model of the environment[1, 2, 3, 4, 5]. The world model reduces random explorations by predicting the outcomes of different actions, leading to a fast convergence speed during online fine-tuning after minimal interactions [6]. Moreover, it significantly improves action safety in high-dimensional tasks such as vision-based operations due to its prediction ability[7, 8]. Consequently, We consider utilizing model-based methods to tackle safety-critical problems in real-world deployment.
@@
Offline RL has been widely studied within the academic community, using pre-collected datasets to learn policies without exploration[9]. At the offline training stage, safety constraints will be incorporated into policy learning, which means learning to be safe other than learning safely. In this way, offline trained policy is promising to avoid constraint violations during the online deployment. However, offline dataset cannot fully simulate complicated real-world environments and will cause out-of-distribution (OOD) issues in the case of unseen data. Such generalization tasks are significant challenges for offline learning.
@@
In this work, we propose developing an offline-to-online approach to address the generalization issue with safety guarantees. We consider the problem of pretraining an offline policy and a world model, and then fine-tuning them safely during online interactions. However, such a strategy usually causes degradation of policy performance and a lack of safety after fine-tuning[10]. The main reason lies in too many mixed constraints (both soft and hard), such as safety constraints and behavior regularization, making it difficult to ensure both optimality and safety. Meanwhile, the world model will introduce prediction errors due to distribution shift during the fine-tuning phase, leading to increasing safety risks. Therefore, we seek to unify and handle mixed constraints, and use the offline safe policy as a guidance for online correction. Based on this key insight, we proposed FOSP (Fine-tuning Offline Safe Policy through World Models), a method that enhances offline safe training and bridges model-based offline training with online fine-tuning.
@@
To evaluate our method, we design different experiments in the simulation including offline training and online fine-tuning. It shows robust performance in both offline and online fine-tuning tasks in Safety-Gymnasium[11] and achieves nearly zero-cost. We also deploy FOSP in a real-world Franka manipulator for trajectory planning tasks. It turns out that our method performs well across tasks with different safety regions and can be transferred to unseen scenarios through few-shot fine-tuning.
@@
We summarize our contributions as follows: (1) To the best of our knowledge, FOSP is the first approach that extends model-based RL to safe offline-to-online areas. (2) It handles the trade-off between the performance and constraints satisfaction across offline and online training phases on various vision-only tasks. (3) In real-world deployment, it does not need sim-to-real transfer and can be safely fine-tuned in unseen safety-constrained scenarios. (4) To inspire further research in this area, we will publicly release our code and dataset.
Figure 1: Fine-tuning offline safe policy through world models. We propose an approach to offline pretrain and online fine-tuning the world model. We also consider the safe constraints and generalize it to unseen safe regions.
@@ 2 Related Works
Offline RL, which trains a policy by leveraging a large amount of existing data, is valuable to real-world scenarios[12, 13]. Prior offline RL works focus on addressing the distribution shift and value overestimation problem[14, 15, 16, 17, 18, 19]. Recently, in-sample methods have demonstrated their robust performance by avoiding out-of-distribution (OOD) state-action pairs[20, 21, 22, 23, 10, 24]. Unlike training solely on offline data, offline-to-online RL involves further fine-tuning the offline policy online. And many offline RL algorithms have extended to online fine-tuning[20, 10]. But these methods have shown that the more effective the offline RL method is, the worse online fine-tuning performance is[25]. Some previous works have explored how to effectively fine-tune from offline pretrained policy, including balanced sampling in replay buffer[26], policy calibrating[27, 28], and parameter transferring[29, 30]. However, existing solutions are restricted by policy transitions or the need to estimate the density of the online policy distribution[31]. PEX[32] is an approach that bridges offline and online algorithms and is adaptive for fine-tuning different online methods with offline policy guidance. As it is shown as an effective approach in offline-to-online training, we extend it to model-based algorithms that not only require fine-tuning the policy but also the model.
@@
Safe RL focuses on optimizing objectives with different constraints[33]. Many methods are based on policy search algorithms that ensure adherence to near-constraints iteratively[34] and Lagrangian methods that convert it to an unconstrained problem to solve[35, 36, 37]. Model-based RL methods utilize world models for action selection so that they ensure safety by background planning. Some recent works are trying to address challenges remaining in vision-input tasks[38, 39, 3]. LAMDBA[7] is a method extending DreamerV1[1] to safe RL by using augmented Lagrangian. Safe-SLAC[40] incorporating the Lagrangian method into SLAC[41] improves the computation complexity of LAMBDA. SafeDreamer[8] is the latest work in safe model-based reinforcement learning. It enhances safety in online or background planning with DreamerV3[5], achieving state-of-the-art (SOTA) performance. However, prior works only aim to address issues in the online setting, which cannot avoid costs incurring during exploration in the real world. By contrast, some safe offline RL can learn policies from offline datasets by using OOD action detection[42, 43, 44], DICE-based theory[45], and sequence modeling[46, 47]. Nonetheless, they only utilize soft constraints which sometimes can not effectively decrease constraint violations. Some works consider the hard constraints by using control theory methods such as control barrier function (CBF)[48] and HJ reachability[49, 50, 51, 52]. Our work draws from these works but unifies all the constraints with model-based background planning.
@@ 3 Preliminaries
Our work extends DreamerV3[5], a powerful and robust baseline model-based RL using world model and actor-critic. It has the ability to model high-dimensional observations of the environment and increase the data efficiency of RL. The world model learns compressed image input representations and predicts future states and rewards sequences for decision-making. It is implemented as a Recurrent State-Space Model (RSSM)[3]. First, the model maps the high-dimensional observations $x_{t}$ to stochastic states $z_{t}$. And deterministic latent states $h_{t}$ are predicted by $z_{t}$ and actions $a_{t}$. By concatenating both of them, we get the model state $s_{t}=\{h_{t},z_{t}\}$. Finally, we facilitate $s_{t}$ to predict the next observations and rewards. Due to the safety considered in the environment, we add a cost decoder to the original model. All components in the world model are as follows:
@@
$\mathcal{L}^{\text{model}}(\theta)=\operatorname*{\mathbb{E}}_{\tau\sim\mathcal{D}}\Big{[}\sum_{t=1}^{T}-\ln p_{\theta}({\bm{x}}_{t}\mid{\bm{s}}_{t})-\ln p_{\theta}({\bm{r}}_{t}\mid{\bm{s}}_{t})-\ln p_{\theta}({\bm{c}}_{t}\mid{\bm{s}}_{t})+$
@@
The stop-gradient operator makes the representations more predictable in the imagination training.
@@
Safe RL is frequently formulated as a Constrained Markov Decision Process (CMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathbb{P},\mathcal{R},\mathcal{C},\mu,\gamma)$[33]. $\mathcal{S}$ and $\mathcal{A}$ denote the state and action space. $\mathbb{P}(s^{\prime}|s,a)$ represents the transition probability from $s^{\prime}$ to $s$ under action $a$. $\mathcal{R}$ is reward space that maps $\mathcal{S}\times\mathcal{A}$ to $\mathbb{R}$. And $\mathcal{C}$ is a cost function set containing cost functions $c:\mathcal{S}\times\mathcal{A}\rightarrow[0,C_{max}]$. $\mu(\cdot):\mathcal{S}\rightarrow[0,1]$ is the initial state distribution and $\gamma\in(0,1)$ is the discount factor. Given a policy distribution $\pi_{\psi}$, we define the performance and constraints as follows:
$J^{\mathcal{R}}(\pi_{\psi})=\operatorname*{\mathbb{E}}_{{\bm{a}}_{t}\sim\pi,{\bm{s}}_{t+1}\sim p_{\theta},{\bm{s}}_{0}\sim\rho}\left[\sum_{t=0}^{T}r_{t}\big{\rvert}{\bm{s}}_{0}\right],$ (2)
$J^{\mathcal{C}}_{i}(\pi_{\psi})=\operatorname*{\mathbb{E}}_{{\bm{a}}_{t}\sim\pi,{\bm{s}}_{t+1}\sim p_{\theta},{\bm{s}}_{0}\sim\rho}\left[\sum_{t=0}^{T}c_{t}^{i}\big{\rvert}{\bm{s}}_{0}\right]\leq d^{i},\;\forall i\in\{1,\dots,C\},$ (3)
@@ 4 Methods
In this section, we introduce our framework (Figure 1) for fine-tuning safe actions with offline pretrained world models. It alleviates the distribution shift error during offline training and balances the trade-off between performance and violations via the reachability estimation function. In the online fine-tuning phase, using offline-trained policy as guidance helps mitigate the performance drop and further improves the performance of the algorithm. In section 4.1, to address the distribution shift in offline training, we use the in-sample optimization approach to train critics. Next, we introduce an offline-to-online approach called model-based policy expansion to further improve online performance in section 4.2. The reachability guidance for safety consideration is provided in section 4.3.
@@ 4.1 In-sample Optimization for Offline Training
Value overestimation is an extensive challenge in offline training due to the distribution shift between the state-action pair from the dataset and generated by the learned policy. Moreover, model-based offline RL algorithms not only learn a policy but also a dynamic model $p_{\theta}$ by the offline dataset. It makes this challenge more severe that the latent dynamic estimation and return prediction are also affected. Therefore, we introduce the in-sample optimization to address this issue.
@@
First, we consider a sequence of data $\mathcal{B}=({\bm{x}}_{1:T},{\bm{a}}_{1:T},{\bm{r}}_{1:T},{\bm{c}}_{1:T})$ sampled from the offline dataset. Then, the model generates latent states ${\bm{s}}_{1:T}\sim q_{\theta}({\bm{s}}_{1:T}|{\bm{h}}_{1:T},{\bm{x}}_{1:T})$ and uses them as initial states to generate rollouts:
$\hat{{\bm{a}}}_{1:T}^{h}\sim\pi_{\psi}({\bm{a}}|\hat{{\bm{s}}}_{1:T}^{h-1}),\ \hat{{\bm{s}}}_{1:T}^{h+1}\sim p_{\theta}({\bm{s}}|\hat{{\bm{a}}}_{1:T}^{h},\hat{{\bm{s}}}_{1:T}^{h}),\ \hat{{\bm{r}}}_{1:T}^{h}\sim p_{\theta}({\bm{r}}|\hat{{\bm{s}}}_{1:T}^{h}),\ \hat{{\bm{c}}}_{1:T}^{h}\sim p_{\theta}({\bm{c}}|\hat{{\bm{s}}}_{1:T}^{h}).$ (5)
@@
We use $Q_{\phi}({\bm{s}},{\bm{a}})$ as the critic network and $\hat{Q}_{\phi}(s,a)$ as the target network. Similar to Implicit Q-learning (IQL)[24], we only use in-sample actions as calculated components of $\lambda$-return. Specifically, a large target value estimated by $Q_{\phi}$ might not represent an existing action leading to a good state due to the out-of-sample actions $\pi({\bm{s}})$. Thus, a separate value network $V_{\varphi}({\bm{s}})=\operatorname*{\mathbb{E}}_{{\bm{a}}\sim\pi(\cdot|{\bm{s}})}[Q_{\phi}({\bm{s}},{\bm{a}})]$ only conditioned on states is proposed to mitigate the out-of-distribution issue. It approaches the action distribution expectation of Q-value and can be optimized by expectile regression:
$\mathcal{L}_{V}(\varphi)=|\tau-\mathds{1}_{\{Q_{\phi}({\bm{s}}_{t}^{j},{\bm{a}}_{t}^{j})-V_{\varphi}({\bm{s}}_{t}^{j})<0\}}|(Q_{\phi}({\bm{s}}_{t}^{j},{\bm{a}}_{t}^{j})-V_{\varphi}({\bm{s}}_{t}^{j}))^{2},$ (6)
where $\tau\in(0,1)$ is a constant. Note that the action ${\bm{a}}$ will be sampled from the dataset according to the loss. Hence, we use rollouts in (5) to estimate returns by TD($\lambda$) (see Appendix.B). The objective can replace the TD-target that was originally used to optimize the Q-value. However, only updating the critic depending on model rollout is not enough, while the real state-action pairs from the offline dataset can supervise the critic to additionally learn in-distribution information. So the critic distribution is optimized on both model data and offline data:
$\mathcal{L}_{Q}^{1}(\phi)=-\frac{1}{HT}\sum_{t=1}^{T}\sum_{h=1}^{H}\log{\text{Pr}_{Q_{\phi}}(R_{t}^{h}|{\bm{s}}_{t}^{h})},\;\;\;\mathcal{L}_{Q}^{2}(\phi)=-\frac{1}{T}\sum_{t=1}^{T}\log{\text{Pr}_{Q_{\phi}}({\bm{r}}_{t}+\gamma V_{\varphi}({\bm{s}}_{t+1})|{\bm{s}}_{t})}$ (7)

### v2

@@ Abstract
Offline Safe Reinforcement Learning (RL) seeks to address safety constraints by learning from static datasets and restricting exploration. However, these approaches heavily rely on the dataset and struggle to generalize to unseen scenarios safely. In this paper, we aim to improve safety during the deployment of vision-based robotic tasks through online fine-tuning an offline pretrained policy. To facilitate effective fine-tuning, we introduce model-based RL, which is known for its data efficiency. Specifically, our method employs in-sample optimization to improve offline training efficiency while incorporating reachability guidance to ensure safety. After obtaining an offline safe policy, a safe policy expansion approach is leveraged for online fine-tuning. The performance of our method is validated on simulation benchmarks with five vision-only tasks and through real-world robot deployment using limited data. It demonstrates that our approach significantly improves the generalization of offline policies to unseen safety-constrained scenarios. To the best of our knowledge, this is the first work to explore offline-to-online RL for safe generalization tasks. The videos are available at https://sunlighted.github.io/fosp_web/.
@@ 1 Introduction
Offline Reinforcement Learning (RL) has been widely studied within the academic community (Wu et al., 2019; Fujimoto et al., 2019; Kostrikov et al., 2021; Kumar et al., 2020), using pre-collected datasets to learn policies without exploration (Levine et al., 2020). At the offline training stage, safety constraints can be incorporated into policy learning, which focuses on learning to be safe rather than learning safely. In this way, offline trained policy is promising to avoid constraint violations during online deployment. However, offline datasets cannot fully simulate complicated real-world environments and will cause out-of-distribution (OOD) issues in the case of unseen data (Kumar et al., 2020). Such generalization tasks are significant challenges for offline learning and safe learning.
This paper aims to explore the design of a practical safe RL algorithm that can both leverage offline data and quickly adapt to novel environments while maintaining safety properties. A straightforward way is to train a safe RL policy on offline data. However, the offline trained policy is prone to distribution shift issues, making it hard to apply to unseen environments due to severe safety constraint violations. Hence, a safe offline-to-online training mechanism is necessary to improve generalization. We consider utilizing model-based methods to complete generalization as quickly as possible. Model-based RL has proven to be a powerful tool to enhance sample efficiency by constructing a world model of the environment (Hafner et al., 2019a; Cang et al., 2021; Hafner et al., 2019b; 2020; 2023). The world model reduces random explorations by predicting the outcomes of different actions, leading to a fast convergence speed during online fine-tuning after minimal interactions (Feng et al., 2023). Moreover, it significantly improves action safety in high-dimensional tasks such as vision-based operations due to its prediction ability (As et al., 2022; Huang et al., 2023). Thus, we leverage its advantages to address the safe generalization problem in real-world deployment.
Figure 1: Fine-tuning offline safe policy through world models. We propose a framework for offline pretraining and online fine-tuning the world model. We first pretrain the agent by the offline dataset and rollouts generated from world models. The grey section depicts the architecture of the world model: it first encodes an image observation into its latent state $s_{0}$, then, for each latent state, generates an action using the policy, as well as predicts the reward, cost, and next state. In the offline-to-online phase, we employ policy expansion to initialize a new policy for online fine-tuning. The pretrained Q-value is leveraged to construct a softmax probability distribution. Then, we select an action by this distribution for the agent to safely interact with the real world, generalizing it to novel tasks.
Figure 2: Safety insurance in FOSP. We enable a safe policy to predict the probability of constraint violations in the future. It can maintain persistent safety in the feasible set and reenter the feasible set as soon as possible when in the infeasible set.
@@
We consider a framework of pretraining an offline safe policy with a world model and then directly fine-tuning it during online interactions. However, such a strategy usually causes degradation of policy performance and a lack of safety after fine-tuning (Nair et al., 2020). The main reason lies in too many mixed constraints (both soft and hard), such as safety constraints and behavior regularization, making it difficult to ensure both optimality and safety. Meanwhile, the world model will introduce prediction errors due to the distribution shift during the fine-tuning on unseen tasks, leading to increasing safety risks. Therefore, we seek to handle mixed constraints and use the offline safe policy as guidance for online correction. Based on this key insight, we propose FOSP (Fine-tuning Offline Safe Policy through World Models), which enhances offline safe training and bridges model-based offline training with online fine-tuning without constraints violations, as shown in Figure 1.
Taking advantage of SafeDreamer (Huang et al., 2023), a safe version of DreamerV3 (Hafner et al., 2023), we employ world models to optimize both offline and online RL, significantly improving data utilization efficiency. Specifically, during offline pretraining, our approach leverages in-sample optimization to update the Q-value conservatively in order to deal with the value overestimation. We also consider safe constraints with a feasibility-guided method, while simultaneously using the reachability estimation function to address the mixed constraints involved in this problem, as illustrated in Figure 2. Moreover, we adopt the safe policy expansion (Zhang et al., 2023) to bridge offline and online algorithms, avoiding performance drops during the initial stage and suboptimality with nearly zero violations. In this way, FOSP balances the trade-off between optimal performance and constraint violations, allowing for safe fine-tuning on generalization tasks.
@@
To evaluate our method, we design experiments across numerous Safety-Gymnasium (Ji et al., 2023) tasks and real-world robotic arms including offline training and online fine-tuning. For simulation tasks, experimental results show robust performance in both offline and online fine-tuning phases and achieve nearly zero cost, outperforming prior RL algorithms. Furthermore, we deploy FOSP in real-world experiments, utilizing a Franka manipulator to perform trajectory planning tasks. It turns out that our method’s ability across tasks with different safety regions can be transferred to unseen safety-critical scenarios through few-shot fine-tuning.
@@
We summarize our contributions as follows: (1) To the best of our knowledge, FOSP is the first approach that tackles safe generalization tasks by offline-to-online RL. (2) It handles the trade-off between the performance and constraints satisfaction across offline and online training phases on various vision-only tasks. (3) In real-world deployment, it does not need sim-to-real transfer and can be safely fine-tuned in unseen safety-constrained scenarios. (4) It can solve offline-to-online safe RL adaptation by only a few trials while maintaining safety.
@@ 2 Related Works
Offline RL, which trains a policy by leveraging a large amount of existing data, is valuable to real-world scenarios (Wu et al., 2019; Fujimoto et al., 2019). Prior offline RL works focus on addressing the distribution shift and value overestimation problem (Yu et al., 2020; 2021; Rafailov et al., 2021; Kumar et al., 2020; Kidambi et al., 2020; Fujimoto & Gu, 2021). Recently, in-sample methods have demonstrated their robust performance by avoiding out-of-distribution (OOD) state-action pairs (Kostrikov et al., 2021; Xu et al., 2023; Garg et al., 2023; Peng et al., 2019; Nair et al., 2020; He et al., 2024; Yang et al., 2024). Unlike training solely on offline data, offline-to-online RL involves further fine-tuning the offline policy online. And many offline RL algorithms have extended to online fine-tuning (Kostrikov et al., 2021; Nair et al., 2020). But these methods have shown that the more effective the offline RL method is, the worse online fine-tuning performance is (Xiao et al., 2023). Some previous works have explored how to effectively fine-tune from offline pretrained policy, including balanced sampling in replay buffer (Lee et al., 2022b), policy calibrating (Nakamoto et al., 2024; Rudner et al., 2021), and parameter transferring (Xie et al., 2021; Rajeswaran et al., 2017). However, existing solutions are restricted by policy transitions or the need to estimate the density of the online policy distribution (Zhao et al., 2022). PEX (Zhang et al., 2023) is an approach that bridges offline and online algorithms and is adaptive for fine-tuning different online methods with offline policy guidance. We extend it to safe algorithms that not only improve the performance but also ensure safety, for it is shown as an effective approach in offline-to-online training.
@@
Safe RL focuses on optimizing objectives with different constraints (Altman, 2021). Many methods are based on policy search algorithms that ensure adherence to near-constraints iteratively (Achiam et al., 2017) and Lagrangian methods that convert it to an unconstrained problem to solve (Chow et al., 2018; Ding et al., 2020; Tessler et al., 2018). Model-based RL methods utilize world models for action selection so that they ensure safety by background planning. Some recent works are trying to address challenges remaining in vision-input tasks (Sikchi et al., 2022; Hansen et al., 2022; Hafner et al., 2019b). LAMDBA (As et al., 2022) is a method extending DreamerV1 (Hafner et al., 2019a) to safe RL by using augmented Lagrangian. Safe-SLAC (Hogewind et al., 2022) incorporating the Lagrangian method into SLAC (Lee et al., 2020) improves the computation complexity of LAMBDA. SafeDreamer (Huang et al., 2023) is the latest work in safe model-based reinforcement learning. It enhances safety in online or background planning with DreamerV3 (Hafner et al., 2023), achieving state-of-the-art (SOTA) performance. However, prior works only aim to address issues in the online setting, which cannot avoid costs incurred during exploration in the real world. By contrast, some safe offline RL can learn policies from offline datasets by using OOD action detection (Cao et al., 2024; Xu et al., 2022; Li et al., 2022), DICE-based theory (Lee et al., 2022a), and sequence modeling (Lin et al., 2023; Liu et al., 2023). Nonetheless, they only utilize soft constraints which sometimes can not effectively decrease constraint violations. Some works consider the hard constraints by using control theory methods such as control barrier function (CBF) (Choi et al., 2020) and HJ reachability (Zheng et al., 2023; Yu et al., 2023; Ganai et al., 2024; Fisac et al., 2019). Our work draws from these works and unifies all the constraints with model-based background planning, continuing the CMDP paradigm that satisfies safety constraints in expectation.
@@ 3 Preliminaries
Our work extends DreamerV3 (Hafner et al., 2023), a powerful and robust baseline model-based RL using world model and actor-critic. It has the ability to model high-dimensional observations of the environment and increase the data efficiency of RL. The world model learns compressed image input representations and predicts future states and rewards sequences for decision-making. It is implemented as a Recurrent State-Space Model (RSSM) (Hafner et al., 2019b). First, the model maps the high-dimensional observations ${\bm{x}}_{t}$ to stochastic states ${\bm{z}}_{t}$. And deterministic latent states ${\bm{h}}_{t}$ are predicted by ${\bm{z}}_{t}$ and actions ${\bm{a}}_{t}$. By concatenating both of them, we get the model state ${\bm{s}}_{t}=\{{\bm{h}}_{t},{\bm{z}}_{t}\}$. Finally, we facilitate ${\bm{s}}_{t}$ to predict the next observations and rewards. Due to the safety considerations in the environment, we add a cost decoder to the original model. All components in the world model are as follows:
@@
$\mathcal{L}^{\text{model}}(\theta)=\mathbb{E}_{\tau\sim\mathcal{D}}\Big{[}\sum_{t=1}^{T}-\ln p_{\theta}({\bm{x}}_{t}\mid{\bm{s}}_{t})-\ln p_{\theta}({\bm{r}}_{t}\mid{\bm{s}}_{t})-\ln p_{\theta}({\bm{c}}_{t}\mid{\bm{s}}_{t})+$
@@
The stop-gradient operator $sg(\cdot)$ makes the representations more predictable in the imagination training.
@@
Safe RL is frequently formulated as a Constrained Markov Decision Process (CMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathbb{P},\mathcal{R},\mathcal{C},\mu,\gamma)$ (Altman, 2021) with discrete time steps $t\in\{0,\ldots,T\}$. $\mathcal{S}$ and $\mathcal{A}$ denote the state and action space. $\mathbb{P}({\bm{s}}^{\prime}|{\bm{s}},{\bm{a}})$ represents the transition probability from ${\bm{s}}^{\prime}$ to ${\bm{s}}$ under action ${\bm{a}}$. $\mathcal{R}$ is reward space that maps $\mathcal{S}\times\mathcal{A}$ to $\mathbb{R}$. And $\mathcal{C}$ is a cost function set containing cost functions ${\bm{c}}:\mathcal{S}\times\mathcal{A}\rightarrow[0,C_{max}]$. $\mu(\cdot):\mathcal{S}\rightarrow[0,1]$ is the initial state distribution and $\gamma\in(0,1)$ is the discount factor. Given a policy distribution $\pi_{\psi}$, the cost state value function $V^{c}$, and cost action-state value function $Q_{c}$ are defined by $V^{c}({\bm{s}})=\mathbb{E}_{\tau\sim\pi_{\psi}}[\sum_{t=0}^{T}\gamma^{t}{\bm{c}}({\bm{s}}_{t},{\bm{a}}_{t})],Q^{c}({\bm{s}},{\bm{a}})=\mathbb{E}_{\tau\sim\pi_{\psi}}[\sum_{t=0}^{T}\gamma^{t}{\bm{c}}({\bm{s}}_{t},{\bm{a}}_{t})]$ like the common MDP. We define the finite-horizon reward function and cost function as follows:
$J^{\mathcal{R}}(\pi_{\psi})=\mathbb{E}_{{\bm{a}}_{t}\sim\pi_{\psi},{\bm{s}}_{t+1}\sim p_{\theta},{\bm{s}}_{0}\sim\mu}\left[\sum_{t=0}^{T}r_{t}\big{\rvert}{\bm{s}}_{0}\right],$ (2)
$J^{\mathcal{C}}_{i}(\pi_{\psi})=\mathbb{E}_{{\bm{a}}_{t}\sim\pi_{\psi},{\bm{s}}_{t+1}\sim p_{\theta},{\bm{s}}_{0}\sim\mu}\left[\sum_{t=0}^{T}c_{t}^{i}\big{\rvert}{\bm{s}}_{0}\right]\leq d^{i},\;\forall i\in\{1,\dots,C\},$ (3)
@@
Reachability Estimation Function
RESPO (Ganai et al., 2024) first introduced the reachability estimation function to capture the probability of constraint violations. Suppose that the set $\mathcal{S}_{v}$ is a constraint violation set defined in the state space $\mathcal{S}$, containing all states that violate the constraints. The reachability estimation function (REF) $u^{\pi}:\mathcal{S}\rightarrow[0,1]$ can be defined as:
$u^{\pi}(s):=\mathbb{E}_{\tau\sim\pi}\Big{[}\max_{{\bm{s}}_{t}\in\tau}\mathds{1}\{({\bm{s}}_{t}|{\bm{s}}_{0},\pi)\in\mathcal{S}_{v}\}\Big{]}$ (5)
The value $\max_{{\bm{s}}_{t}\in\tau}\mathds{1}\{({\bm{s}}_{t}|{\bm{s}}_{0},\pi)\in\mathcal{S}_{v}\}$ is defined on specific trajectory. It equals 1 if existing violations and 0 otherwise.
@@ 4 Methods
In this section, we introduce FOSP for fine-tuning safe actions with offline pretrained world models. According to Figure 1, we first train an offline safe policy and subsequently fine-tune it online. In Section 4.1, we address the Q-value overestimation in offline learning by utilizing in-sample optimization to train the critic network $Q_{\phi}({\bm{s}},{\bm{a}})$. In Section 4.2, to balance the trade-off between performance and constraint violations during offline training, we employ the reachability estimation function to train the safe actor $\pi_{\psi}({\bm{s}})$. In Section 4.3, the policy is online fine-tuned by safe policy expansion, which helps mitigate the performance drop and enables further improvements. The whole algorithm can be found in Appendix C.
@@ 4.1 In-sample Optimization for Offline Training
Value overestimation is an extensive challenge in offline RL due to the distribution shift between the state-action pairs from the dataset and those generated by the learned policy. This issue becomes even more pronounced in model-based offline RL algorithms, where both the policy and a dynamics model $p_{\theta}$ are learned from the offline data. It not only amplifies value overestimation but also causes additional complications in estimating the latent dynamics and predicting future returns, further impacting the overall performance. To mitigate these challenges, we propose the in-sample optimization as a solution.
@@
First, we consider a sequence of data $\mathcal{B}=({\bm{x}}_{1:T},{\bm{a}}_{1:T},{\bm{r}}_{1:T},{\bm{c}}_{1:T})$ sampled from the offline dataset. Then, the model generates latent states ${\bm{s}}^{0}_{1:T}\sim q_{\theta}({\bm{s}}_{1:T}|{\bm{h}}_{1:T},{\bm{x}}_{1:T})$, which are used as initial states $\hat{{\bm{s}}}^{0}_{1:T}$ to generate rollouts:
$\hat{{\bm{a}}}_{1:T}^{h}\sim\pi_{\psi}({\bm{a}}|\hat{{\bm{s}}}_{1:T}^{h}),\ \hat{{\bm{s}}}_{1:T}^{h+1}\sim p_{\theta}({\bm{s}}|\hat{{\bm{a}}}_{1:T}^{h},\hat{{\bm{s}}}_{1:T}^{h}),\ \hat{{\bm{r}}}_{1:T}^{h}\sim p_{\theta}({\bm{r}}|\hat{{\bm{s}}}_{1:T}^{h}),\ \hat{{\bm{c}}}_{1:T}^{h}\sim p_{\theta}({\bm{c}}|\hat{{\bm{s}}}_{1:T}^{h}),$ (6)
@@
where $h$ represents the horizon of the generated sequence. We use $Q_{\phi}({\bm{s}},{\bm{a}})$ as the critic network. Motivated by Implicit Q-learning (IQL) (Kostrikov et al., 2021), we only use in-sample actions as calculated components of $\lambda$-return. Due to out-of-distribution actions from $\pi_{\psi}({\bm{s}})$, a high Q-value from $Q_{\phi}$ doesn’t always mean the action will lead the agent to a desirable state. Thus, a separate value network $V_{\varphi}({\bm{s}})=\mathbb{E}_{{\bm{a}}\sim\pi_{\psi}(\cdot|{\bm{s}})}[Q_{\phi}({\bm{s}},{\bm{a}})]$ only conditioned on states is proposed to mitigate the out-of-distribution issue. It approaches the action distribution expectation of Q-value and can be optimized by expectile regression:
$\mathcal{L}_{V}(\varphi)=\mathbb{E}_{\tau\sim\mathcal{D},\pi_{\psi},p_{\theta}}\Big{[}|\kappa-\mathds{1}\{Q_{\phi}(\hat{{\bm{s}}}_{t}^{j},\hat{{\bm{a}}}_{t}^{j})-V_{\varphi}(\hat{{\bm{s}}}_{t}^{j})<0\}|(Q_{\phi}(\hat{{\bm{s}}}_{t}^{j},\hat{{\bm{a}}}_{t}^{j})-V_{\varphi}(\hat{{\bm{s}}}_{t}^{j}))^{2}\Big{]},$ (7)
where $\kappa\in(0,1)$ is a constant. We use rollouts in equation 6 to estimate returns by TD($\lambda$) and update the critic by equation 9.
$R_{t}^{H}=V_{\varphi}(\hat{{\bm{s}}}_{t}^{H}),\;R_{t}^{h}=\hat{{\bm{r}}}_{t}^{h}+\gamma((1-\lambda)V_{\varphi}(\hat{{\bm{s}}}_{t}^{h})+\lambda R_{t}^{h+1}),$ (8)
$\mathcal{L}_{Q}^{1}(\phi)=-\frac{1}{HT}\mathbb{E}_{\tau\sim\mathcal{D},\pi_{\psi},p_{\theta}}\Big{[}\sum_{t=1}^{T}\sum_{h=1}^{H}(Q_{\phi}(\hat{{\bm{s}}}^{h}_{t},\hat{{\bm{a}}}^{h}_{t})-R^{h}_{t})^{2}\Big{]},$ (9)
where $H$ is the horizon of the model and $h\leq H$. It is noted that the $\lambda$-return in equation 9 is estimated solely through the generated rewards. However, during the early stages of offline training, the world model may not be able to generate accurate rewards. The real rewards from the offline dataset can better guide the critic’s learning process. Therefore, we add equation 10 as a regularization term to the critic loss to improve learning efficiency:
$\mathcal{L}_{Q}^{2}(\phi)=-\frac{1}{T}\mathbb{E}_{\tau\sim\mathcal{D},\pi_{\psi},p_{\theta}}\Big{[}\sum_{t=1}^{T}[Q_{\phi}({\bm{s}}_{t}^{0},{\bm{a}}_{t})-({\bm{r}}_{t}+\gamma V_{\varphi}({\bm{s}}_{t+1}^{0}))]^{2}\Big{]},$ (10)